OpenStudy (anonymous):

please check this answer \[\lim_{x \rightarrow +\infty} \frac{ 2^{4 } - 3x}{ x ^{4}+1}\] x=10000 \[\frac{ 2*(10000)^{4} - 3*10000 }{ (10000)^{4} + 1}\] = \[- \infty\]

OpenStudy (anonymous):

It is true

OpenStudy (anonymous):

It is like \[ -\ frac 3 {x^3{ \]

OpenStudy (ash2326):

@Muskan please check again

OpenStudy (anonymous):

divide both numerator and denominator with x^4. then apply the limits.

OpenStudy (anonymous):

\[ -\frac 3 {x^3} \]

OpenStudy (anonymous):

When x is near Inifinty, the fraction is near - Infinity

OpenStudy (anonymous):

\[\frac{ 2 }{ 1 }\]

OpenStudy (anonymous):

muskan its\[\lim_{x \rightarrow +\infty} \frac{ 2x^{4 } - 3x}{ x ^{4}+1}\]ha?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

i think , it should be 2 then instead of -infinity.

OpenStudy (anonymous):

when \(x\rightarrow \infty\)\[x^4>>x>>1\]so u can neglect \(3x\) in the num and \(1\) in denum In comparison to \(x^4\) so\[\lim_{x \rightarrow +\infty} \frac{ 2x^{4 } - 3x}{ x ^{4}+1}=\lim_{x \rightarrow +\infty} \frac{ 2x^{4 } }{ x ^{4}}\]

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

The original question was \[ \lim_{x \rightarrow +\infty} \frac{ 2^{4 } - 3x}{ x ^{4}+1} \] and this limit is zero